OpenStudy (anonymous):

If(a-bcos(y))(a+bcos(x))=a^(2)-b^(2),show that (dy)/(dx)=(√(a^(2)+b^(2)))/(a+bcos(x))?

OpenStudy (primeralph):

product rule.

OpenStudy (loser66):

@e.mccormick what is it?

OpenStudy (loser66):

if you say so, ok, show me steps. I have no idea to come up with the problem.

OpenStudy (loser66):

come here friend, we do it.

OpenStudy (loser66):

got it, i got until \[y'= \frac{absinx -b^2cosy}{absiny+abcosx}\]

OpenStudy (loser66):

give me more time, still not get the answer yet. still not see the link . haizzz

OpenStudy (loser66):

no, i find out multiply everything is much more easier than our stuff... hahaha. redo, friend,

OpenStudy (loser66):

LHS = \[a^2 +abcosx-abcosy-b^2cosycosx= cosy(-ab-b^2cosx)+a^2+abcosx\] from this I can isolate y , right? the whole thing = a^2 + b^2 Isolate cosy and then take implicit derivative, yeah...

OpenStudy (loser66):

no ok? you think so? :( , let me try.

OpenStudy (loser66):

I hand off, too tired. friend, I give up.... too long. and no outlet. hehe. I don't have your patience